A bank issues a $1,440 loan to a borrower, who will repay the loan with a probability of 86%. In this case, the competitive interest rate the bank can afford to offer is

Group of answer choices

21%

18.8%

16.3%

14.5%
Answer #1

Suppose bank charges a rate of X%. Then bank will be receiving 1440 + 14.40X in case the loan is repaid. The expected amount of loan bank will be receiving is E(L) = 0.86*(1440 + 14.40X) + 0.14*0 and this should be enough to cover the loan of 1440. Hence we have

0.86*(1440 + 14.40X) = 1440

1238.40 + 12.384X = 1440

X = (1440 – 1238.40)/12.384 = 16.3%

Select 16.3%